# Keyword-Aware ASR Error Augmentation for Robust Dialogue State Tracking

### v1

@@ 1 Introduction
Figure 1: Illustration of EPA process.
@@
However, DST models face significant challenges in spoken dialogue environments, where user utterances are converted into text by automatic speech recognition (ASR). Notably, Soltau et al. (2022) observed a drastic reduction in model accuracy from 41.6% to 23.6% in such environments [^1] Joint goal accuracy result of base size model.. This decline is primarily due to ASR errors, which frequently misrecognize named entities—a key target in DST Nechaev et al. (2021).
To address ASR inaccuracies, data augmentation has emerged as a viable, cost-efficient strategy. Existing text augmentation methods, such as word swapping Wei and Zou (2019) and back translation Sennrich et al. (2015), do not maintain audio similarity with the original text, leading to discrepancies with ASR error patterns. To bridge this gap, Sharma et al. (2020) and Jacqmin et al. (2023) synthesized audio from text with text-to-speech (TTS) model Shen et al. (2018) and processed it through ASR, while Hrinchuk et al. (2020) and Zhang et al. (2021) employed translation model structure to introduce ASR-like errors directly into texts.
Despite these advancements, prior methods often fail to provide sufficient error patterns for DST model training. Accurately identifying key terms is vital for DST performance; thus, models need to be trained on a broad spectrum of ASR-errored keywords. Unfortunately, many current strategies do not ensure that errors are positioned within critical keywords, often generating trivial examples by altering non-essential words such as random words Wei and Zou (2019) or sentence structure Sennrich et al. (2015). This oversight results in sub-optimal DST performance against ASR errors.
@@
In our method, we utilize in-context Brown et al. (2020) learning to guide the augmentation process. By providing examples that highlight the key parts, we enable the model to position errors in keywords. Surprisingly, without collecting extensive user voice datasets, a manually created small text dataset (fewer than 10 samples) can generate a wide variety of ASR-errored keywords. This simplifies the error generation process. Additionally, before keyword-focused augmentation, we augmented the text to incorporate utterance-level overall ASR errors, introducing diverse errors from open-source audio datasets. This two-step approach effectively harnesses LLMs for both overall utterance-level and targeted keyword-level phonetic augmentation.
In our experiments, EPA achieved a diversity in targeted keywords (95.4%) and high variance in overall text (1.81x) while maintaining audio similarity (91.5%). During practical DST evaluations, EPA significantly enhanced the robustness of DST models, increasing accuracy from 45.76% to 51.12%. Our analysis indicates that these improvements are primarily due to keyword augmentation, which significantly reduces errors in ASR-errored values.
@@ 2.1 Notation
Before detailing each step, we first clarify the notation. Dialogue context from turn 1 to $t$ is denoted as $D_{t}$=$\{(s_{1},u_{1}),...,(s_{t},u_{t})\}$ where $s$ denotes for system and $u$ for user utterance. DST model predicts the dialogue state (also called belief state) $B_{t}$ given $D_{t}$. $B_{t}$ is composed with slot $sl$ and value $v$ pairs, denoted as $B_{t}$ = $\{({sl}_{1},v_{1}),...,({sl}_{J},v_{J})\}$ , where ${sl}_{j}$ and $v_{j}$ is $j$-th slot name and value. $J$ is the total number of slots.
@@
3 Original Tuesday, going to bailey’s crossroads please.
+EPA Tuesday, going to baley’s crossroads, peas.
@@ 4 Original Hi! Could you find me a train to loris on thursday?
Table 1: Examples of ASR errors from EPA. We have highlighted utterance level overall errors in yellow and keyword-specific errors in blue.
2.2 Step 1: ASR Error for Overall Utterance
Next, we augment the overall utterance level of $u$ by guiding the LLM using examples retrieved from $DB$ based on audio similarity (see Figure  1, Step 1). To calculate audio similarity, we converted the characters to phonemes using the International Phonetic Alphabet (IPA) 222 https://pypi.org/project/eng-to-ipa/ and then computed the similarity between the phonemized $u$ and $DB$. After selecting the $k$ examples, we concatenated the instructions, examples, and $u$ into a single prompt and provided it to the LLM. This process resulted in the overall ASR-errored user utterance $\dot{u}$. Concretely, $\dot{u}$ can be obtained by
$\dot{u_{t}}=LLM(Inst_{1}\oplus(g_{1},e_{1})\cdots(g_{k},e_{k})\oplus u_{t}).$ (1)
2.3 Step 2: ASR Error for Keywords
@@
Table 2: Comparison of various augmentation methods in enhancing the robustness of DST models across different ASR environments. All results were averaged over three seeds for better consistency.
While Step 1 introduces general ASR errors into ${u}$, it does not ensure that these errors enhance the diversity of keywords. To create a more effectivetraining dataset, we generated keyword-focused ASR errors (see Figure 1, Step 2).Here, we manually created several examples from the original DST dataset by explicitly highlighting the keywords in the clean text using the <hl> tag ($\tilde{g}$) and concatenated them with the ASR-errored versions ($\tilde{e}$). We set the $v$ in turn $t$ as keywords. We applied the same highlighting to the inference target $\dot{u}$ and then concatenated the instructions, examples and highlighted $\dot{u}$ into a single prompt, as in Step 1. We perform the in-context learning with this prompt to generate the augmented utterance $\ddot{u}$, containing overall ASR and keyword-specific errors. Explicitly, we get $\ddot{u}$ by
$\ddot{u_{t}}=LLM(Inst_{2}\oplus(\tilde{g}_{0},\tilde{e}_{0})\cdots(\tilde{g}_{k},\tilde{e}_{k})\oplus\dot{u}_{t}).$ (2)
We set the example numbers as 3 in both Step1 and Step2.
After generating the ASR-errored examples, we trained a DST model using both the EPA-generated data and the original seed data. Specifically, we fine-tuned T5-base Roberts et al. (2019) models for DST tasks, with detailed training procedures outlined in the Appendix B.
3 Augmentation Results of EPA
We first assessed the quality of ASR errors simulated by the EPA method to guarantee they closely resembled genuine ASR errors. Table 1 demonstrates that with overall utterance and keyword error augmentation, EPA effectively generates sentences that align with our objectives: maintaining phonetic similarity to the original text while confirming errors in keywords.
4 Experiments
@@ 4.1 Experimental Setup
Dataset. The DSTC11 dataset Soltau et al. (2022), an audio version of MultiWOZ 2.1 Eric et al. (2019), comprises 8,000 dialogues for training, 1,000 for validation, and 1,000 for testing. To enhance generalization, we conducted experiments across four distinct ASR environments, characterized by word error rate (WER) [^3] Based on LibriSpeech.test-clean Radford et al. (2022). and noise levels: (1) a low accuracy ASR model (WER > 0.03), (2) a café and traffic noised audio, (3) a paraphrased setting where users naturally paraphrased the transcriptions, and (4) a high accuracy ASR model (WER < 0.03) enviroment.
@@
LLMs. we primarily used GPT-3.5 Ouyang et al. (2022). Additionally, as this is the first attempt to augment pronunciation-related errors with LLMs, we also included Llama2-7B Touvron et al. (2023) and OPT-6.7B Zhang et al. (2022).
Compared methods. We compared our method with two established approaches: text-based augmentations, AEDA Karimi et al. (2021), EDA Wei and Zou (2019), and Back Translation (BT) Sennrich et al. (2015), and audio-aware augmentation methods, using synthesized audio (TTS-ASR) and translation model structure (ASR-translation). Lastly, we included Olisia Jacqmin et al. (2023), the top-ranked method in the DSTC11 competition (Appendix C).
@@ 4.2 Robustness Improvement through EPA
The results in Table 2 demonstrate the effectiveness of EPA in enhancing robustness to ASR errors in the DST task. Remarkably, EPA outperformed existing text-based and audio-based augmentation methods across all environments, showing substantial improvement in N-Acc with gains of over 4%, compared to the modest 1-2% improvements of other methods. This trend is particularly pronounced in challenging environments, such as low-accuracy ASR and noisy audio conditions, and is also observed in LLaMA-7B and OPT 6.7B. The ablation study revealed that keyword error augmentation consistently boosted performance across all LLMs, capitalizing on its utility to optimize DST performance against ASR errors.
@@ 4.3 Error Analysis
High-acc ASR (WER<0.03)
Baseline
$\triangledown$0%
$\triangledown$0%
$\triangledown$0%
EPA w/o Key-aug
$\triangledown$1.57%
$\triangledown$3.34%
$\triangledown$6.52%
EPA
$\triangledown$7.16%
$\triangledown$2.43%
$\triangledown$0.75%
Table 3: Ablation study with error analysis. Wrong indicates the model predicts incorrect values, Ignore refers to the model ignored mentioned slots, and Spurious denotes the model predicting values for unmentioned slots. Actual error numbers are in parentheses.
To examine the effects of keyword augmentation in more detail, we conducted ablation studies and analyzed error reduction rates by categorizing errors into three types. Table 3 presents the percentage reduction in error rates compared to the baseline. The results demonstrate that EPA is effective in "Wrong" and "Ignore" error types, and keyword augmentation highly contributed to this improvement by decreasing the error rate from 5.29% to 8.19%. Interestingly, while keyword augmentation led to substantial reductions in the "Wrong" type error, it increased the "Spurious" error type. This suggests that the model may become overly cautious, occasionally overestimating the presence of named entities.
4.4 Qualitative Assessment of EPA Method
Method Uniq. Words NE.chg [ % ] Audio Sim.[%]
Baseline 1 - -
AEDA Karimi et al. (2021) 1.00$\times$ 44.29 91.57
EDA Wei and Zou (2019) 0.86$\times$ 70.03 61.14
BT Sennrich et al. (2015) 1.21$\times$ 73.46 77.17
TTS-ASR 1.01$\times$ 38.84 98.93
Translating 0.84$\times$ 39.59 94.07
EPA 1.81$\times$ 95.47 91.57
   w/o Keyword Err. 1.57$\times$ 68.81 93.14
Table 4: Assessment of EPA dataset quality across three key aspects (Unique word increased rate(Uniq. Words, Named entity changed rate (NE.chg), and audio similarity (Audio Sim.).
To evaluate the word diversity and audio similarity of the augmented text, we performed a qualitative assessment in Table 4. This assessment includes metrics such as the unique word increase rate, named entity change rate, and audio similarity  [^4] Normalized phoneme edit distance is used for audio similarity calculating.. The results demonstrate that EPA achieves remarkable diversity in unique words (1.81x) and the highest rate of named entity augmentation (95.47%), while maintaining considerable audio similarity (91.57%). Furthermore, keyword augmentation significantly contributes to the improvement, particularly in the named entity change rate (from 68.81% to 91.57%), ensuring that the errors exhibit sufficiently diverse patterns in keywords.

### v2

# Speak & Spell: LLM-Driven Controllable Phonetic Error Augmentation for Robust Dialogue State Tracking

@@ Abstract
[…]
Speak & Spell: LLM-Driven Controllable Phonetic Error Augmentation for Robust Dialogue State Tracking
Jihyun Lee1, Solee Im3thanks: This work was conducted at POSTECH., Wonjun Lee2, Gary Geunbae Lee1,2 1Graduate School of Artificial Intelligence, POSTECH, Republic of Korea 2Department of Computer Science and Engineering, POSTECH, Republic of Korea 3KT {jihyunlee, lee1jun, gblee}@postech.ac.kr {solee.im}@kt.com
@@ 1 Introduction
Figure 1: Illustration of the EPA process. Step 1 introduces overall ASR-like distortions using in-context examples of original–errored pairs, while Step 2 applies keyword-specific phonetic corruptions guided by highlighted slot spans. Combining both steps generates realistic and diverse ASR-style errors for robust DST training.
@@
However, DST models face significant challenges in spoken dialogue environments, where user utterances are converted into text by automatic speech recognition (ASR) Pal et al. (2020); Kim et al. (2021); Yoon et al. (2023). Notably, Soltau et al. (2022) observed a drastic reduction in model accuracy from 41.6% to 23.6% in such environments. This decline is primarily due to ASR errors, which frequently misrecognize named entities—a key target in DST Nechaev et al. (2021).
To address ASR inaccuracies, data augmentation has emerged as a viable, cost-efficient strategy. Existing text augmentation methods, such as word swapping Wei and Zou (2019) and back translation Sennrich et al. (2015), do not maintain audio similarity with the original text, leading to discrepancies with ASR error patterns. To bridge this gap, Sharma et al. (2020) and Jacqmin et al. (2023) synthesized audio from text with text-to-speech (TTS) model Shen et al. (2018) and processed it through ASR, while Hrinchuk et al. (2020) and Zhang et al. (2021) employed translation model structure to introduce ASR-like errors directly into texts. Huang and Chen (2020) further leveraged word confusion network (WCN) representations for data augmentation to generate acoustically consistent errors.
Despite these advancements, prior methods often fail to provide sufficient error for DST model training. Accurately identifying key terms is vital for DST performance; thus, models need to be trained on a broad spectrum of ASR-errored keywords. Unfortunately, many current strategies do not ensure that errors are positioned within critical keywords, often generating trivial examples by altering non-essential words such as random words Wei and Zou (2019) or sentence structure Sennrich et al. (2015). This oversight results in sub-optimal performance against ASR errors.
@@
In our method, we utilize in-context learning Brown et al. (2020) with phonetically similar examples to introduce general ASR errors and devise a highlighting method to explicitly localize the error to a target span. Surprisingly, without requiring extensive domain-specific user speech data, a publicly available audio dataset and a small set of in-context examples (fewer than 10 samples) are sufficient to generate a wide variety of ASR-errored keywords for DST. This significantly simplifies the error generation process.
In the experiment, to reflect diverse real-world conditions, we evaluated EPA under four ASR environments: a low-accuracy ASR system, noisy audio with café and traffic background, a paraphrased input setting where users naturally rephrased transcriptions, and a high-accuracy ASR system. In these experiments, EPA significantly improved model robustness, increasing accuracy from 45.76% to 51.12% with high keyword diversity (95.4%), surpassing the previous best-performing model. Our analysis suggests that this improvement is primarily driven by keyword-level augmentation, which effectively mitigates errors in ASR-affected values.
@@ 2.1 Notation
Before detailing each step, we first clarify the notation. Dialogue context from turn 1 to $t$ is denoted as $D_{t}$$\{(s_{1},u_{1}),...,(s_{t},u_{t})\}$ where $s$ denotes for system and $u$ for user utterance. DST model predicts the dialogue state (also called belief state) $B_{t}$ given $D_{t}$. $B_{t}$ is composed with slot $sl$ and value $v$ pairs, denoted as $B_{t}$ = $\{({sl}_{1},v_{1}),...,({sl}_{J},v_{J})\}$ , where ${sl}_{j}$ and $v_{j}$ is $j$-th slot name and value. $J$ is the total number of slots.
2.2 Step 1: ASR Error for Overall Utterance
Next, we inject errors into $u$ by prompting the LLM with in-context examples. We retrieved ($g$, $e$) pairs from the database ($DB$) based on phonetic similarity between $u$ and $g$ (Figure 1, Step 1). To compute phonetic similarity, we converted the characters of both $u$ and $g$ into phonemes using the International Phonetic Alphabet (IPA), and calculated similarity using a frequency-based retrieval  [^1] We used BM25Robertson et al. (2009), a retrieval model based on term frequency. While neural retrievers (e.g., DPRKarpukhin et al. (2020)) could be applied, we opted for a frequency-based method, as neural models tend to capture semantic similarity.. After selecting the top-$k$ ($g$, $e$) pairs, we concatenated the instruction, in-context examples, and $u$ into a single prompt and provided it to the LLM. This process results in the overall ASR-errored user utterance, denoted as $\dot{u}$. Concretely, $\dot{u}$ can be obtained by
$\dot{u_{t}}=LLM(Inst_{1}\oplus(g_{1},e_{1})\cdots(g_{k},e_{k})\oplus u_{t})$ (1)
where $\oplus$ denotes concatenation, and we set $k=3$ throughout our experiments. Retrieved examples are provided in Appendix A.3.
2.3 Step 2: ASR Error for Keywords
While Step 1 introduces general ASR-style errors into $u$, it does not ensure sufficient error diversity in keyword tokens. To construct a more effective training dataset, we explicitly generate keyword-focused ASR errors in Step 2 (Figure 1). In this step, we highlight the keywords in $\dot{u}$ using the <hl> tag and instruct the LLM to inject errors specifically within the highlighted spans. For the DST task, we treat dialogue state values ($v$) as keywords, although the definition of a keyword may vary depending on the task. To facilitate this process, we provide a few examples that illustrate how values within <hl> tags are intended to be modified during augmentation. Given these instructions and examples, the LLM generates an augmented utterance $\ddot{u}$ that includes both general and keyword-specific ASR errors. Formally, we obtain $\ddot{u}_{t}$ as follows:
$\ddot{u_{t}}=LLM(Inst_{2}\oplus({g}_{0},{e}_{0})\cdots({g}_{k},{e}_{k})\oplus\dot{u}_{t}).$ (2)
The used prompts are provided in Appendix A.1.
2.4 Examples of EPA
@@
1 Original Tuesday, going to bailey’s crossroads please.
+EPA Tuesday, going to baley’s crossroads, peas.
@@ 4 Original Hi! Could you find me a train to loris on thursday?
Table 1: Examples of ASR errors from EPA.
Table 1 shows examples of ASR errors generated by EPA. We have highlighted utterance level overall errors in yellow and keyword-specific errors in orange. For instance, in Row 1, the model introduces a keyword-level error (bailey’s → baley’s) as well as an additional phonetically plausible insertion (peas), simulating realistic ASR noise. Further examples can be found in Appendix 9.
3 Experiments
@@
Table 2: Comparison of various augmentation methods in enhancing the robustness of DST models across different ASR environments. In the feature columns, “Aud.” indicates the use of audio-based augmentation, “Utt-aug.” denotes utterance-level text augmentation, “Key-aug.” refers to keyword-specific augmentation, and “LLM” indicates whether large language models were used for error generation. All results were averaged over three seeds for better consistency.
@@ 3.1 Experimental Setup
Dataset. The DSTC11 dataset Soltau et al. (2022), an audio version of MultiWOZ 2.1 Eric et al. (2019), comprises 8,000 dialogues for training, 1,000 for validation, and 1,000 for testing. To enhance generalization, we conducted experiments across four distinct ASR environments, characterized by Word Error Rate (WER) and noise levels: (1) a low accuracy ASR model (WER > 0.03), (2) a café and traffic noised audio, (3) a paraphrased setting where users naturally paraphrased the transcriptions, and (4) a high accuracy ASR model.
@@
Compared methods. We compared our method with two established approaches: text-based augmentations, AEDA Karimi et al. (2021), EDA Wei and Zou (2019), and Back Translation (BT) Sennrich et al. (2015), and audio-aware augmentation methods, using synthesized audio (TTS-ASR) and translation model structure (ASR-translation). Lastly, we included Olisia Jacqmin et al. (2023), the top-ranked method in the DSTC11 competition.
Models. For performing EPA, we used diverse types of LLMs, including GPT-3.5Ouyang et al. (2022), LLaMA2-7BTouvron et al. (2023) and OPT-6.7BZhang et al. (2022). For the DST task, we fine-tuned a T5-baseRoberts et al. (2019) model. Further details about the experimental settings are provided in Appendix B.
@@ 3.2 Robustness Improvement through EPA
EPA improves robustness.
The results in Table 2 shows the effectiveness of EPA in robustness to ASR errors. Remarkably, EPA outperformed existing text-based and audio-based augmentation, showing substantial improvement in JGA and named entity accuracy. It also surpassed the previous best-performing model, Olisia, particularly in challenging environments.
Effectiveness of keyword-specific error.
In Table 2, we present an ablation study to evaluate the effectiveness of keyword-level augmentation. We found that adding keyword-specific ASR errors improved DST performance across all environments and was particularly helpful in enhancing the robustness of named entity accuracy. Additional experiments, including generalization to other backbones and tasks, as well as statistical significance analysis, are provided in Appendix C.
3.3 Comparison with Error Correction Models
Method Weak-ASR Noised-ASR Strong-ASR
JGA N-acc JGA N-acc JGA N-acc
Baseline 29.88 45.76 29.70 46.77 34.87 52.07
+ Correction model 1 23.30 41.50 23.96 42.86 26.68 46.14
+ Correction model 2 30.14 46.79 29.82 47.68 34.62 52.00
+ EPA (with GPT, ours) 32.39 51.12 32.24 52.70 36.61 55.87
Table 3: Comparison with ASR error correction models under different ASR conditions.
In Table 3, we compare our method with two ASR error correction approaches. For Error Correction Model 1, We fine-tuned a pre-trained on 300 hours of ASR-errored and gold transcription pairs (same as in Section 2.2) in a seq2seq manner, minimizing $L=-\sum_{i=1}^{N}\log P(g_{i}\mid e_{i})$, where $g_{i}$ and $e_{i}$ denote the gold and errored transcriptions, respectively. Training used early stopping (patience=3) with batch size 16. For Error Correction Model 2, we adopted an off-the-shelf222urlhttps://huggingface.co/oliverguhr/spelling-correction-english-base text correction model to revise typos and transcription errors. Each model corrected the test transcriptions from the weak, noised, and strong ASR systems before DST evaluation. As in the table, Correction model 1 underperforms the baseline, often altering named entities unnecessarily (e.g., "Grifon" $\rightarrow$ "Bristol"), likely due to limited training diversity. Correction model 2 yields minor improvements, while our EPA achieves the highest robustness across all ASR conditions.
3.4 Qualitative Assessment of EPA Method
Method
Uniq.
Words
NE.chg
Pronoun
Sim.[%]
Baseline 1 - -
[…]
Table 4: Assessment of EPA dataset quality: Unique word increased rate, Named entity changed rate (NE.chg), and pronunciation similarity.
Automatic evaluation.
Although Table 2 confirms EPA’s effectiveness, it remains unclear whether the LLM-generated augmentations truly reflect diverse, keyword-focused ASR-style errors. To this end, we perform the quality analysis based on three metrics (Table 4): the unique word increase rate, named entity change rate, and pronunciation similarity with original sentence. The results show that EPA achieves remarkable diversity in unique words (1.81×) and the highest named entity change rate (95.47%), while maintaining high pronunciation similarity (91.57%). Notably, keyword-level augmentation plays a key role in enhancing named entity variability, increasing the change rate from 68.81% to 95.47%.
Comparison with Authentic ASR Errors.
Figure 2: Distribution of edit distance. The x-axis represents edit distances, and the y-axis represents the corresponding ratio.
Method Text Dist.($\downarrow$) Phoneme Dist.($\downarrow$)
ASR-L ASR-B ASR-L ASR-B
TTS-ASR (Whisper-B) 0.030 0.048 0.056 0.061
TTS-ASR (Whisper-S) 0.030 0.048 0.070 0.077
ASR trans. (T5-base) 0.025 0.039 0.104 0.116
EPA (Llama2 7B) 0.218 0.123 0.204 0.256
EPA (OPT 6.7B) 0.115 0.106 0.071 0.091
EPA (GPT 3.5) 0.033 0.010 0.009 0.007
Table 5: Distribution distance (JSD) between Whisper Large/Base model and simulation dataset. (ASR‑L=Whisper Large, ASR‑B=Whisper Base).
In this analysis, we explored the similarity between simulated ASR errors and authentic ASR errors from the perspective of edit distance. Specifically, we examined the distribution of edit distances in simulated data and in errors produced by Whisper large/base models, considering both character- and phoneme-level representations (Figure 2). To quantify the distributional differences, we computed the Jensen-Shannon Divergence (JSD), a symmetric variant of the Kullback-Leibler divergence (Table 5).
Our experiments yielded several noteworthy findings. Notably, the LLM-simulated errors closely matched the distribution of real ASR errors, particularly at the phoneme level. This suggests that LLMs can effectively capture pronunciation-level variations and generate realistic ASR-style distortions. In contrast, errors produced by TTS-based synthesis or ASR-translation models exhibited larger divergence from real ASR patterns, likely due to their tendency to generate more diverse but less phonetically grounded outputs.
Human evaluation.
To further verify the quality of our EPA method, we conducted a human evaluation using 100 sentence pairs, each consisting of an original sentence and its augmented counterpart, with two human evaluators. Participants rated how likely the change resembled an ASR error on a 4-point Likert scale, where 1 indicated "not like an ASR error" and 4 indicated "clearly an ASR error." The average rating was 3.22 with moderate inter-rater agreement (Gwet’s AC2 Gwet (2008) = 0.590), suggesting that most EPA-generated edits were perceived as realistic ASR errors. Details on the evaluation metric and human evaluation are provided in Appendix D.
@@ 3.5 Error Analysis
We additionally analyze the impact of keyword augmentation by examining how it influences specific error types in DST predictions. Table 6 presents the percentage reduction in error rates compared to the baseline. The results demonstrate that EPA is effective in "Wrong" and "Ignore" error types, and keyword augmentation highly contributed to this improvement by decreasing the error rate from 5.29% to 8.19%. Interestingly, while keyword augmentation led to substantial reductions in "Wrong" errors, it also caused a slight increase in "Spurious" errors. This may be because the model, after repeatedly seeing phonetic noise around slot values, becomes overly sensitive and starts hallucinating unmentioned slots. A potential mitigation is to introduce an additional loss term for slot presence prediction Heck et al. (2020); Kim et al. (2019) , helping the model better distinguish between mentioned and unmentioned slots.
@@
Table 6: Ablation study with error analysis. Wrong indicates the model predicts incorrect values, Ignore refers to ignored mentioned slots, and Spurious denotes predicting values for unmentioned slots. Actual error numbers are in parentheses.